OpenStudy (anonymous):

StArAnGeL: 5x^4 y^2 (2x^5 y^6) / 20x^3 y^5

OpenStudy (anonymous):

\[\frac{5x^4 y^2 (2x^5 y^6) }{ 20x^3 y^5 }\] Can you try ?

OpenStudy (anonymous):

uh.....er....uh....

OpenStudy (anonymous):

Okay \[\huge \frac{5 \times 2 \times x^4y^2x^5y^6}{20 x^3y^5}\]

OpenStudy (anonymous):

Now ? :)

OpenStudy (anonymous):

Do the same way as we did the previous problem

OpenStudy (anonymous):

LOL! Kay...one sec,plz...

OpenStudy (anonymous):

Note: \[\huge x^m \times x^n =x^{m+n}\]\[\huge x^m \div x^n = x^{m-n}\]

OpenStudy (anonymous):

Is it fraction 1/2 x^6/y^3 ??

OpenStudy (anonymous):

If you meant this one You are absolutely right :) \[\huge \frac{1}{2} x^6y^3 \]

OpenStudy (anonymous):

Why isnt it over y^3 though???

OpenStudy (anonymous):

I get it.THANK YOU!!!!!!! You are my new favorite :) your better than my teacher! Your on often right? (Please say yes)

OpenStudy (anonymous):

I get it.THANK YOU!!!!!!! You are my new favorite :) your better than my teacher! Your on often right? (Please say yes)

OpenStudy (anonymous):

Okay lets consider y we have \[\huge \frac{y^2 \times y^6}{y^5} \] Which is equal to\[\huge \frac{y^{2+6}}{y^5}= \frac{y^8}{y^5}=y^{8-5}=y^3\]

OpenStudy (anonymous):

Glad to hear that :D i will try to help you whenever i am on

OpenStudy (anonymous):

!!! :D :D :D